OpenStudy (anonymous):

sandwich theorem?? Can you explain it clearly.I dont ge it at all.Thank you so much

OpenStudy (anonymous):

you mean about limits?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I say you should have a look at the first example on wiki page. It can't be stated more clearly than that http://en.wikipedia.org/wiki/Squeeze_theorem

OpenStudy (anonymous):

nice picture there too illustrating the idea

OpenStudy (anonymous):

Math noton on that pagis confusing me

OpenStudy (anonymous):

Suppose you are a plane flying into a SANDWICH. Well, there is bread below you, and bread above you. And it's coming together. Well, there's only ONE place you can end up and that's among the bologna. So if the BOTTOM bread is at height 800 meters and the TOP bread is at height 800.004 meters, then both YOU and the BOLOGNA (sounds like bolony) must be at about 800.002 meters.

OpenStudy (anonymous):

so if f(x) is the bottom bread and g(x) is the top bread, and p(x) is the plane. Then because you are between the breads, we know f(x) < p(x) < g(x). Now as x goes to infinity, f(x) approaches 800 and g(x) also approaches 800. Then we know that as x approaches infinity, 800 <= p(x) <= 800. Well there is only one number between 800 and 800, and that's 800! So p(x) = 800.

OpenStudy (anonymous):

merit badge please

OpenStudy (anonymous):

thank you thank you!

OpenStudy (anonymous):

LOL @mareoraft , you are asking for it for quite some time. Thought I would give you one :)

OpenStudy (anonymous):

can you solve an example function using this propery

OpenStudy (anonymous):

it is most appreciated

OpenStudy (anonymous):

Sure, give us the example problem.

OpenStudy (anonymous):

lim x approaches 0 [x^2sin(1/x)]

OpenStudy (anonymous):

See the first example at wikipedia. It is the same problem!

OpenStudy (anonymous):

I read it and didnt understand it there

OpenStudy (anonymous):

Tell me what exactly did you not understand ther ?

OpenStudy (anonymous):

*there

OpenStudy (anonymous):

You just need to know that -1 < sin x < 1

OpenStudy (anonymous):

Multiply x^2 to all the 3 -x^2 < x^2 sin x < x^2

OpenStudy (anonymous):

Apply limit x tends to 0 you get 0< (limit x--> 0) x^2 sin x <0 Therefore (limit x--> 0) x^2 sin x = 0

OpenStudy (anonymous):

why is - <sin 1/x <1

OpenStudy (anonymous):

Read about sin function

OpenStudy (anonymous):

LIM as x approaches 0 OF [x^2sin(1/x)] -1 <= sin(1/x) <= 1 -x^2 <= x^2*sin(1/x) <= x^2 but now we let x go to 0, and we get... 0 <= x^2*sin(1/x) <= 0 so the answer is 0.

OpenStudy (anonymous):

Here http://en.wikipedia.org/wiki/Sine

OpenStudy (anonymous):

I know -1<sinx<1 but why -1<sin 1/x<1

OpenStudy (anonymous):

it's -1 <= sin(y) <= 1 NOT -1 < sin(y) < 1

OpenStudy (anonymous):

I knoq just didnt know how to put the notation here

OpenStudy (anonymous):

it doesn't matter what you put in the sin(). nomatter what you put in, its still bounded by -1 and 1. -1 <= sin(15u^7) <= 1